Alvin paddled for 5 hours with a 5 km/hour current to reach a campsite. The return trip against the same current took 10 hours. Find the speed of Alvin’s canoe without a current.
Answer by lwsshak3(11628)

Alvin paddled for 5 hours with a 5 km/hour current to reach a campsite. The return trip against the same current took 10 hours. Find the speed of Alvin’s canoe without a current.
let x=speed of Alvin’s canoe without a current
(x+5)=speed of canoe with the current
(x-5)=speed of canoe against the current
travel time*speed=distance
5(x+5)=10(x-5)
5x+25=10x-50
5x=75
x=15
speed of Alvin’s canoe without a current=15km/hr